OpenStudy (anonymous):

What is the value of the expression below when x = –2? 5 + 3(x – 7) – (x + 1) A. –5 B. –7 C. –21 D. –23

OpenStudy (anonymous):

I hope this helps: f(x) = 5 + 3(x - 7) - (x + 1) f(-2) = 5 + 3( -2 - 7 ) - ( -2+1) = 5 - 6 - 21 + 2 -1 = - 21

OpenStudy (anonymous):

so the answer is c